# General Phrase Debiaser: Debiasing Masked Language Models at a Multi-Token Level

### 25 Dec 2023 v2

@@ 3.2 Evaluation Result And Analysis
We run General Phrase Debiaser in both career field and discipline field at the same time. The main results of experiments shown in Table 1 demonstrate that our method is capable of reducing model biases, lowering original average scores of BERT, ALBERT, and DistilBERT in the six SEAT tests from 0.35, 0.72, and 0.79, respectively, to 0.12, 0.16, and 0.52. Furthermore, compared to other approaches in the entire benchmark, including those relying on manual datasets or generating data automatically, General Phrase Debiaser shows the state-of-the-art debiasing performance. We find the global superiority of our model is derived from three aspects through analysis:

### 25 Jan 2024 v3

@@ 3.2 Evaluation Result And Analysis
We run General Phrase Debiaser in both career field and discipline field at the same time. The effect size score of the SEAT[16] benchmark we report in Table 1 measures the association between two sets of target concepts and two sets of attributes. It is obtained by calculating the normalized distance between a set of attribute sentence vectors and a set of concept sentence vectors output by the model, and the closer this distance is to 0, the less biased the model is. The result demonstrates that our method is capable of reducing model biases, lowering original average scores of BERT, ALBERT, and DistilBERT in the six SEAT tests from 0.35, 0.72, and 0.79, respectively, to 0.12, 0.16, and 0.52. Furthermore, compared to other approaches in the entire benchmark, including those relying on manual datasets or generating data automatically, General Phrase Debiaser shows the state-of-the-art debiasing performance. We find the global superiority of our model is derived from three aspects through analysis:
@@ 4 Conclusion
Acknowledgements
This work was supported by Grant 2020YFB1005400 from the National Key R&D Program of China.